Evaluate the ∫(t³+3t/t²+1)dt


1
Expert's answer
2021-02-24T12:30:34-0500

Solution:

"\mathrm{Apply\:the\:Sum\:Rule}:\quad \int f\left(x\right)\pm g\left(x\right)dx=\int f\left(x\right)dx\pm \int g\left(x\right)dx"


"\int t^3dt=\dfrac{t^4}{4}"


"\int \dfrac{3t}{t^2}dt=3ln|t|"


"\int 1dt=t"


"\int \:t^3dt+\int \frac{3t}{t^2}dt+\int \:1dt=\dfrac{t^4}{4}+3ln|t|+t+C"


Answer:

"\dfrac{t^4}{4}+3ln|t|+t+C"